# Computationally-Efficient Linear Periodically Time-Variant Digital PLL Modeling Using Conversion Matrices and Uncorrelated Upsampling

### v2

@@ III-B DCO-Rate Transfer Function Analysis
Figure 6: DCO-rate LPTV model for FPEC PLL.
After converting all noise sources to the DCO rate, noise transfer functions are derived utilizing Fig. 6. The divider is modeled as multiplying the output phase by a $N$-periodic pulse, $w_{R,1}[n]/N$, that samples the output phase every $N$ samples:
$w_{R,1}[n]=\begin{cases}1,&n=kN\\ {0},&{\text{else}}\end{cases}$ (20)
for $k\in\{0,1,2,\ldots\}$. The time-varying loop gain implements (16) by multiplying a constant $NK_{P0}/P$ with a $N$-periodic window $w_{R,P}[n]$:
$w_{R,P}[n]=\begin{cases}1,&kN\leq n\leq kN+P-1\\ {0},&{\text{else.}}\end{cases}$ (21)
for $k\in\{0,1,2,\ldots\}$. Multiplying with (20) or (21) falls under the case depicted in Fig. 3, and the associated conversion matrices,$\mathbf{W}_{R,1}$ and $\mathbf{W}_{R,P}$, can be calculated using (15).
@@
    $\mathbf{H}_{ZOH,N}$: conversion matrix of $N$-length zero-order hold $H_{ZOH,N}(z)=\frac{1-z^{-N}}{1-z^{-1}}$.
@@
    $\mathbf{K}_{I}$: conversion matrix of the integral path transfer function $\frac{K_{I}}{1-z^{-1}}$.
@@
    $\mathbf{H}_{DCO}$: conversion matrix of the DCO transfer function $H_{DCO}=\frac{K_{DCO}}{f_{DCO}}\frac{z^{-1}}{1-z^{-1}}$. $K_{DCO}$ is the DCO gain with unit Hz/LSB.
The conversion matrices for $\phi_{n,ref,up}[n]$, $q_{n,DSM,up}[n]$, and $\phi_{n,DCO}[n]$ to the output $\phi_{n,PLL}[n]$ are
$\begin{split}\mathbf{H}_{ref,tot}&=(\mathbf{I}-\mathbf{L})\backslash\mathbf{P}_{ref}\\ \mathbf{H}_{DSM,tot}&=(\mathbf{I}-\mathbf{L})\backslash\mathbf{P}_{DSM}\\ \mathbf{H}_{DCO,tot}&=(\mathbf{I}-\mathbf{L})\backslash\mathbf{P}_{DCO}\end{split}$ (22)
@@
$\mathbf{L}=-K_{PD}\mathbf{H}_{DCO}\underbrace{\left(\frac{NK_{P0}}{P}\overbrace{\mathbf{W}_{R,P}}^{\text{FPEC}}\mathbf{H}_{ZOH,N}+\mathbf{K}_{I}\right)}_{\text{Loop Filter}}\overbrace{\frac{\mathbf{W}_{R,1}}{N}}^{\text{Divider}}$ (23)
@@
$\begin{split}\mathbf{P}_{ref}&=K_{PD}\mathbf{H}_{DCO}\left(\frac{NK_{P0}}{P}\mathbf{W}_{R,P}\mathbf{H}_{ZOH,N}+\mathbf{K}_{I}\right)\\ \mathbf{P}_{DSM}&=\mathbf{H}_{DCO}\mathbf{H}_{ZOH,N}\\ \mathbf{P}_{DCO}&=\mathbf{I}\end{split}$ (24)
@@
$\mathcal{L}(\Omega)=\frac{1}{2}S_{\phi,PLL}(\Omega),$ (25)
in which $S_{\phi,PLL}(\Omega)$ is the SSB PSD of $\phi_{n,PLL}[n]$.
@@
$S_{xx,corr}(\Omega)\leftarrow S_{xx}(\Omega)\uparrow L$triangleright Upsampling by $L$, (17)
Compute $\mathbf{W}_{R,1}$, $\mathbf{W}_{R,P}$, $\mathbf{W}_{U,L}$triangleright Window CM, (15)
for $i=1:\text{length}(\mathbf{\Omega}_{vec})$ do
     $\Omega_{0}\leftarrow\mathbf{\Omega}_{vec}(i)$
     Evaluate $\mathbf{H}_{LTI}(\Omega_{0})$triangleright (9)
     $\mathbf{L}(\Omega_{0})\leftarrow f(\mathbf{W}_{R},\mathbf{H}_{LTI}(\Omega_{0}))$triangleright Loop gain, (23)
     $\mathbf{P}(\Omega_{0})\leftarrow g(\mathbf{W}_{R},\mathbf{H}_{LTI}(\Omega_{0}))$triangleright Path gain, (24)
     $\mathbf{H}(\Omega_{0})\leftarrow(\mathbf{I}-\mathbf{L}(\Omega_{0}))\backslash\mathbf{P}(\Omega_{0})$
     $\mathbf{H}_{U,L}(\Omega_{0})\leftarrow\mathbf{H}(\Omega_{0})\mathbf{W}_{U,L}$triangleright Decorrelation, (19)
     $\mathcal{L}(\Omega_{0})\leftarrow h(\mathbf{H}_{U,L}(\Omega_{0}),S_{xx,corr}(\Omega_{0}))$triangleright (5)
@@
In Algorithm 1, $\mathbf{\Omega}_{vec}$ contains frequency values of interest sampled within $[0,\pi)$, and the algorithm solves a $N\times N$ linear system of equations each time with all LTI transfer functions evaluated at $\Omega_{0}=\mathbf{\Omega}_{vec}(i)$. Compared with simulation-based phase noise extraction, the proposed algorithm computes $\mathcal{L}(\Omega_{0})$ almost instantaneously, delivering accuracy comparable to transient simulations that might be time-consuming.
@@ IV Simulation Results
The proposed algorithm, which calculates the output PN using conversion matrices, is validated with Simulink transient simulations. The Simulink model implements Fig. 6 with parameters shown in Table I. The TDC resolution, $\Delta t_{res}$, is set to 15.16 ps, leading to $K_{PD}=1/(2\pi f_{ref}\Delta t_{res})$ = 300 $\text{rad}^{-1}$ referring to the reference frequency $f_{ref}=35$ MHz. The reference is assumed to have much less phase noise than the TDC, and is thus neglected in the following simulations.
@@
$K_{PD}$ 300 $\text{rad}^{-1}$
$K_{P0}$ 0.4
$K_{I}$ $K_{P0}/32$
$K_{DCO}$ 4 MHz/LSB
$f_{ref}$ 35 MHz
@@ IV-A DCO
With excellent modeling accuracy provided above, the elapsed time of the proposed algorithm is 0.67 s for 16000 frequency domain samples, 21$\times$ shorter than the Simulink simulation time consumption. The efficiency of the algorithm can be further leveraged if only a few, instead of 16000, spot PN needs to be optimized, allowing the designer to quickly explore the design space.
@@
Figure 8: PN and jitter improvement with different FPEC window length $P$.
@@ IV-B DSM
Contrary to the continuous-time analysis predicting that the DSM noise has a high-pass shape (similar to the dashed green and blue lines) in the output PN, the actual low-frequency PN is dominated by aliased DSM noise. The output PN at 1 MHz and the integrated jitter is shown in Fig. 10 with different DSM clock divide ratio, $M$, for $N$ = 18 and $N$ = 16. With smaller $M$, i.e., larger $f_{DSM}$, the quantization noise is pushed to higher frequency, resulting in a general trend of smaller RMS jitter. The 1 MHz spot PN, however, first increases with larger $f_{DSM}$, peaks at $M/2$, then decreases. When the in-band PN is critical in the application, $M$ needs to be carefully selected to avoid excessive noise aliasing. If the above results were to be generated without uncorrelated upsampling, the time consumption will be much larger under certain cases, e.g., 125$\times$ for $N=18$ and $M=5$.
@@ IV-C TDC
The simulated output PN due to TDC quantization noise is shown in Fig. 11 for FPEC window length $P=2$ and in the absence of FPEC ($P=18$). The TDC input-referred quantization noise floor is $\mathcal{L}_{TDC,in}=1/(12)/K_{PD}^{2}/f_{ref}=-135.8$ dBc/Hz. The PLL output PN has a low-pass shape with a low-frequency plateau $25.120\log N=$ 25.1 dB higher than the TDC input-referred noise floor. The colored simulation results matches well with the black lines predicted by the model. In the case of $P=2$, the output jitter due to TDC is 3.43 ps, 1 dB higher than 3.04 ps when $P=18$.
@@
IV-D Total Output PN
Fig. 13 shows the output PN with all noise sources activated: TDC quantization noise, DSM quantization noise, and DCO phase noise. When $P$ reduces from 18 to 2, the in-band PN shows a 2.2 dB improvement, while the out-of-band noise plateau due to DSM quantization noise rises by 1.3 dB. Fig. 14 shows the in-band PN and RMS jitter improvement in finer resolution of $P$. The integrated RMS jitter improves only 0.93 dB when $P$ reduces from 18 to 2, a result of the trade-off between lower DCO noise and higher reference/TDC noise.
@@ IV-D Total Output PN
IV-E Algorithm Complexity
Fig. 15 shows the computation time required for a single frequency-domain sample using the proposed conversion matrix-based algorithm across varying problem dimensions $N$, specifically when $M=N-1$ and $P=2$. This algorithm was evaluated on a laptop equipped with an AMD Ryzen 9 5900HS CPU and 16.0 GB of RAM. For values where $N<150$, computation time exhibits a relationship proportional to $N^{2}$. However, for greater dimensions ($N\geq 200$), the complexity appears to be proportional to $N^{2.5}$. The proposed uncorrelated upsampling approach can reduce computation time by a factor of at least $(N-1)^{2}$. When $N<65$, the computation time for one frequency-domain sample remains under 1 ms, enabling sweeps across the design space to fine-tune PN at selected frequencies with great accuracy.

### v3

@@ III-B DCO-Rate Transfer Function Analysis
Figure 7: DCO-rate LPTV model for the FPEC PLL. The zero-order hold (ZOH) block $H_{\text{ZOH},N}$ is used to keep the value of the feedback signal constant during one reference period $T_{\text{REF}}$, and $H_{\text{ZOH},M}$ holds the upsampled DSM noise constant during $T_{\text{DSM}}$.
After converting all noise sources to the DCO rate, noise transfer functions are derived utilizing Fig. 7. The divider is modeled as multiplying the output phase by a $N$-periodic pulse, $w_{\text{R},1}[n]/N$, that samples the output phase every $N$ samples:
$w_{\text{R},1}[n]=\begin{cases}1,&n=kN\\ {0},&{\text{else}}\end{cases}$ (20)
for $k\in\{0,1,2,\ldots\}$. The time-varying loop gain implements (16) by multiplying a constant $NK_{\text{P0}}/P$ with a $N$-periodic rectangular window $w_{\text{R},P}[n]$:
$w_{\text{R},P}[n]=\begin{cases}1,&kN\leq n\leq kN+P-1\\ {0},&{\text{else.}}\end{cases}$ (21)
for $k\in\{0,1,2,\ldots\}$. Multiplying with (20) or (21) falls under the case depicted in Fig. 3, and the associated conversion matrices,$\mathbf{W}_{\text{R},1}$ and $\mathbf{W}_{\text{R},P}$, can be calculated using (15).
@@
    $\mathbf{H}_{\text{ZOH},N}$: conversion matrix of $N$-length zero-order hold $H_{\text{ZOH},N}(z)=\frac{1-z^{-N}}{1-z^{-1}}$, where $z=e^{j\Omega}$.
@@
    $\mathbf{K}_{\text{I}}$: conversion matrix of the integral path transfer function $\frac{K_{\text{I}}}{1-z^{-1}}$.
@@
    $\mathbf{H}_{\text{DCO}}$: conversion matrix of the DCO transfer function $H_{\text{DCO}}=\frac{K_{\text{DCO}}}{f_{\text{DCO}}}\frac{z^{-1}}{1-z^{-1}}$ [23]. $K_{\text{DCO}}$ is the DCO gain with unit Hz/LSB.
The conversion matrices for $\phi_{\text{n,ref,up}}[n]$, $q_{\text{n,DSM,up}}[n]$, and $\phi_{\text{n,DCO}}[n]$ to the output $\phi_{\text{n,PLL}}[n]$ are
$\begin{split}\mathbf{H}_{\text{REF,tot}}&=(\mathbf{I}-\mathbf{L})^{-1}\mathbf{P}_{\text{REF}}\\ \mathbf{H}_{\text{DSM,tot}}&=(\mathbf{I}-\mathbf{L})^{-1}\mathbf{P}_{\text{DSM}}\\ \mathbf{H}_{\text{DCO,tot}}&=(\mathbf{I}-\mathbf{L})^{-1}\mathbf{P}_{\text{DCO}}\end{split}$ (22)
@@
$\mathbf{L}=-K_{\text{PD}}\mathbf{H}_{\text{DCO}}\underbrace{\left(\frac{NK_{\text{P0}}}{P}\overbrace{\mathbf{W}_{\text{R},P}}^{\text{FPEC}}\mathbf{H}_{\text{ZOH},N}+\mathbf{K}_{\text{I}}\right)}_{\text{Loop Filter}}\overbrace{\frac{\mathbf{W}_{\text{R},1}}{N}}^{\text{Divider}}$ (23)
@@
$\begin{split}\mathbf{P}_{\text{REF}}&=K_{\text{PD}}\mathbf{H}_{\text{DCO}}\left(\frac{NK_{\text{P0}}}{P}\mathbf{W}_{\text{R},P}\mathbf{H}_{\text{ZOH},N}+\mathbf{K}_{\text{I}}\right)\\ \mathbf{P}_{\text{DSM}}&=\mathbf{H}_{\text{DCO}}\mathbf{H}_{\text{ZOH},M}\\ \mathbf{P}_{\text{DCO}}&=\mathbf{I}\end{split}$ (24)
@@
$\mathcal{L}(\Omega)=\frac{1}{2}S_{\phi,\text{PLL}}(\Omega),$ (25)
in which $S_{\phi,\text{PLL}}(\Omega)$ is the SSB PSD of $\phi_{\text{n,PLL}}[n]$.
@@
$S_{yy,\text{corr}}(\Omega)\leftarrow S_{xx}(\Omega)\uparrow L$triangleright Upsampling by $L$, (17)
Compute $\mathbf{W}_{\text{R},1}$, $\mathbf{W}_{\text{R},P}$, $\mathbf{W}_{\text{U},L}$triangleright Window CM, (15)
for $i=1:\text{length}(\mathbf{\Omega}_{\text{vec}})$ do
     $\Omega_{0}\leftarrow\mathbf{\Omega}_{\text{vec}}(i)$
     Evaluate $\mathbf{H}_{\text{LTI}}(\Omega_{0})$triangleright (9)
     $\mathbf{L}(\Omega_{0})\leftarrow f(\mathbf{W}_{R},\mathbf{H}_{\text{LTI}}(\Omega_{0}))$triangleright Loop gain, (23)
     $\mathbf{P}(\Omega_{0})\leftarrow g(\mathbf{W}_{R},\mathbf{H}_{\text{LTI}}(\Omega_{0}))$triangleright Path gain, (24)
     $\mathbf{H}(\Omega_{0})\leftarrow(\mathbf{I}-\mathbf{L}(\Omega_{0}))^{-1}\mathbf{P}(\Omega_{0})$
     $\mathbf{H}_{\text{U},L}(\Omega_{0})\leftarrow\mathbf{H}(\Omega_{0})\mathbf{W}_{\text{U},L}$triangleright Decorrelation, (19)
     $\mathcal{L}(\Omega_{0})\leftarrow h(\mathbf{H}_{\text{U},L}(\Omega_{0}),S_{yy,\text{corr}}(\Omega_{0}))$triangleright (5)
@@
In Algorithm 1, the left arrow $\leftarrow$” represents the assignment of the evaluated expression on the right side to the variable on the left, $\mathbf{\Omega}_{\text{vec}}$ contains frequency values of interest sampled within $[0,\pi)$, and the algorithm solves a $N\times N$ linear system of equations each time with all LTI transfer functions evaluated at $\Omega_{0}=\mathbf{\Omega}_{\text{vec}}(i)$.
@@ IV Simulation Results
The proposed algorithm, which calculates the output PN using conversion matrices, is validated with Simulink transient simulations. The Simulink model implements Fig. 7 with parameters shown in Table I. The TDC resolution, $\Delta t_{\text{res}}$, is set to 15.16 ps, leading to $K_{\text{PD}}=1/(2\pi f_{\text{REF}}\Delta t_{\text{res}})$ = 300 $\text{rad}^{-1}$ referring to the reference frequency $f_{\text{REF}}=35$ MHz. The reference is assumed to have much less phase noise than the TDC, and is thus neglected in the following simulations.
@@
$K_{\text{PD}}$ 300 $\text{rad}^{-1}$
$K_{\text{P0}}$ 0.4
$K_{\text{I}}$ $K_{\text{P0}}/32$
$K_{\text{DCO}}$ 4 MHz/LSB
$f_{\text{REF}}$ 35 MHz
@@ IV-A DCO
Figure 9: PN and jitter improvement with different FPEC window length $P$. RMS jitter is obtained by integrating PN from 10 kHz to $f_{\text{DCO}}/2$.
@@ IV-B DSM
Contrary to the continuous-time analysis predicting that the DSM noise has a high-pass shape (similar to the dashed green and blue lines) in the output PN, the actual low-frequency PN is dominated by aliased DSM noise. The output PN at 1 MHz and the integrated jitter is shown in Fig. 11 with different DSM clock divide ratio, $M$, for $N$ = 18 and $N$ = 16. With smaller $M$, i.e., larger $f_{\text{DSM}}$, the quantization noise is pushed to higher frequency, resulting in a general trend of smaller RMS jitter. The 1 MHz spot PN, however, first increases with larger $f_{\text{DSM}}$, peaks at $M/2$, then decreases. When the in-band PN is critical in the application, $M$ needs to be carefully selected to avoid excessive noise aliasing. If the above results were to be generated without uncorrelated upsampling, the time consumption will be much larger under certain cases, e.g., 125$\times$ for $N=18$ and $M=5$.
@@
Fig. 12 compares the accuracy of the proposed work with [22]. When $N=32$ and $M=5$, [22] generates the same accurate result as this work because $\text{gcd}(M,N)=1$, which falls into the special case 1) mentioned in Section III-A. When $M=2$, $\text{gcd}(M,N)=2$, making [22] predict 2.1 dB lower close-in aliased PN than the simulation, which is the same as this work but without the decorrelation sequence $w_{\text{U},L}[n]$.
Figure 12: Simulated output PN due to DSM quantization noise with $M=2,5$ and $N=P=32$. The proposed model generates correct results regardless of $M$ and $N$, while [22] is only accurate when $\text{gcd}(M,N)=1$.
@@ IV-C TDC
The simulated output PN due to TDC quantization noise is shown in Fig. 13 for FPEC window length $P=2$ and in the absence of FPEC ($P=18$). The TDC input-referred quantization noise floor is $\mathcal{L}_{\text{TDC,in}}=1/(12)/K_{\text{PD}}^{2}/f_{\text{REF}}=-135.8$ dBc/Hz. The PLL output PN has a low-pass shape with a low-frequency plateau $25.120\log N=$ 25.1 dB higher than the TDC input-referred noise floor. The colored simulation results matches well with the black lines predicted by the model. In the case of $P=2$, the output jitter due to TDC is 3.43 ps, 1 dB higher than 3.04 ps when $P=18$.
@@
IV-D Total Output PN
Fig. 15 shows the output PN with all noise sources activated: TDC quantization noise, DSM quantization noise, and DCO phase noise. When $P$ reduces from 18 to 2, the in-band PN shows a 2.2 dB improvement, while the out-of-band noise plateau due to DSM quantization noise rises by 1.3 dB. Fig. 16 shows the in-band PN and RMS jitter improvement in finer resolution of $P$. The integrated RMS jitter improves only 0.93 dB when $P$ reduces from 18 to 2, a result of the trade-off between lower DCO noise and higher reference/TDC noise. When using a mid-rise TDC, the linearized PD gain becomes dependent on the RMS jitter shown at its input. Therefore, the curve of the output RMS jitter vs. $P$ becomes convex while remaining monotonic.
An intuitive explanation of FPEC in the frequency domain is given as follows: Fig. 17 shows the conversion matrix of the loop filter without the integral path for $K_{\text{P0}}=0.4$, $N=4$, and $1\leq P\leq 4$ ($P=4$ disables FPEC). When $P=4$, the conversion matrix of the proportional path $\mathbf{H}$ only provides gain at the baseband (no frequency conversion). When $P=1$, FPEC creates identical copies of the baseband PN at higher frequencies. These upconverted PN, after being filtered by the DCO, alias back to the baseband due to divider sampling, which adds on top of the original PN. This effectively increases the loop gain, helping to suppress DCO noise, as shown in Fig. 8. In other words, FPEC suppresses in-band noise by “seeking help” from the out-of-band gain.
This DPLL can also be understood as a loopback transceiver [43], where the proportional path with FPEC functions as the mixer in the transmitter, the DCO serves as the power amplifier, the divider acts as a mixer-first receiver, and the PD works as an error amplifier in the baseband. The FPEC enables “carrier aggregation” [1], where multiple “channels” shown as colored rectangles in Fig. 17 are used simultaneously to transmit the baseband data (PN), thereby increasing the received data strength.
Figure 17: Intuitive explanation of FPEC in the frequency domain using conversion matrix. The proportional path is composed of $\text{ZOH}_{N}$ and $K_{P}[n]$.
IV-E DSM Fractional Resampling and Delay
Figure 18: Loop filter in Fig. 7 with modified integral path considering the DSM control code fractional resampling and delay.
This subsection revisits the effect of fractional resampling and delay in the integral path introduced by the DSM. The modified loop filter that takes them into account is shown in Fig. 18. Following the integrator, the DSM resampling sequence $w_{M,1}[n]$ with period $M$ is multiplied. An $M^{\text{th}}$-order ZOH is used to keep the value constant within one period, and finally, 2 delays at $f_{\text{DSM}}$ are converted to $z^{-2M}$ at $f_{\text{DCO}}$. The elements in the resampling conversion matrix $\mathbf{H}_{\text{MUL}}$ can be derived from (13) and (15) as follows:
$H_{\text{MUL}(i,j)}=\frac{1}{\text{lcm}(N,M)}W_{0}\left(\frac{2\pi(i-j)}{N}\right),$ (26)
in which
$W_{0}(\Omega)=\sum_{n=0}^{\text{lcm}(N,M)-1}w_{M,1}[n]e^{-jn\Omega}.$ (27)
Fig. 19 compares the results of the simplified loop filter in Fig. 7 and the modified one in Fig. 18 for $K_{\text{P}}/K_{\text{I}}=32$ and $K_{\text{P}}/K_{\text{I}}=2$. In practical cases where $K_{\text{I}}\ll K_{\text{P}}$, it is safe, depending on the required accuracy, to ignore the dynamics in the integral path except for the integrator itself.
Figure 19: Simulated output PN due to DSM quantization noise for $N=18$, $M=5$, and $P=4$. The simplified loop filter in Fig. 7 generates negligible error for $K_{\text{P}}/K_{\text{I}}=32$.
Note that (26) and (27) do not provide an exact solution for fractional resampling as they ignore some of the frequency shifts introduced by the beat frequency between $f_{\text{REF}}$ and $f_{\text{DSM}}$. Fig. 20 shows the simulated output PN due to DSM quantization noise with different $M$. $K_{\text{P}}/K_{\text{I}}$ is set to 2 to not attenuate the error from the integral path. Based on observations, using (26) gives negligible error for all $M\leq N/2$, and the error starts to increase when $M$ approaches $N$ while still being relatively small.
Figure 20: Simulated output PN due to DSM quantization noise for $N=28$, $P=4$, and $M=5,12,22$ considering DSM fractional resampling.
IV-F Cross Verification
Fig. 21 replicates the result of an MDLL modeled in [24, Fig. 14] using the proposed method to demonstrate flexibility. Once the framework is built, the only work left to reproduce the result is rewriting the LTI-like transfer functions (22)-(24) by inspection according to the block diagram shown in [24, Fig. 11].
Fig. 22 explores the validity of the proposed method with reference noise being white and colored (--20 dB/dec). While this work generates correct predictions in both cases, removing the decorrelation sequence generates large error spikes at integer multiples of $f_{\text{REF}}$ for colored input noise. This is a result of violating special case 2) and applying (5) on an $N$-cyclostationary upsampled sequence, similar to Fig. 12. This effect can be avoided by assuming white reference noise and ignoring the frequency shift [24, eq. (30)], [25, eq. (41)], then applying the result to colored reference noise, which ultimately gives negligible error.
Figure 21: Simulated output PN of an MDLL shown in [24] and calculated PN using the proposed model.
Figure 22: Simulated output PN of an MDLL shown in [24] due to colored and white reference noise. The proposed work generates correct predictions under both cases, while removing the decorrelation sequence generates large spikes at integer-multiple of $f_{\text{REF}}$ when the reference noise is colored.